OpenStudy (anonymous):

Evaluate ((3)/(5))^(2)

OpenStudy (jhannybean):

(3/5)^2 or 3/(5)^2?

OpenStudy (anonymous):

\[\left(\frac35\right)^2=\frac{3^2}{5^2}=\frac{9}{25}\]